Consider a reaction with an activation energy of 11090.1 J/mol run at 298.15 K and at 2228 K. How many times larger is the rate constant, k , for the reaction at 2228 K than at 298 K?
Answer #1

Activation energy = 11090.1 J/mol

Temperature T1 = 298.15

T2 = 2228 K

rate constant k1 . k2

ln (k2 / k1) = Ea / R [1 / T1 - 1 / T2]

ln (k2 / k1) = 11090.1 / 8.314 [1/298.15 - 1/2228]

k2 / k1 = 48.2

number of times = 48 times